OpenStudy (savvygrl108):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one.

OpenStudy (savvygrl108):

\[f(x)=\frac{ x-7}{ x+3} and g(x)=\frac{ -3x-7 }{ x-1 }\]

OpenStudy (savvygrl108):

I already have f(g(x)) but I keep struggling on g(f(x))

OpenStudy (savvygrl108):

\[=\frac{ -3\frac{ x-7 }{ x+3 } -7}{ \frac{ x-7 }{ x+3 }-1}\]

OpenStudy (savvygrl108):

That is the set up right?